RC meeting and Run Plan, Thursday November 3rd

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Time:at 3:30pm today

  • Location: Counting house conference room

Useful Links

Meeting Time

  • We have officially changed the meeting time moving forward to 1530. We should be mindful that this gives little time to have Radcon support any plans adjusted during the RC meeting same-day

Current Status

Polarized Target

  • Current Polarization: N/A (Testing over last night, target removed today)
  • Requesting hourly screenshots of Target Strip Charts (and keep an eye on these during shift) - should be removed after configuration change ended
  • Target experts were in overnight to take measurements. All of these measurements are now complete
  • We will not need to return to power permit for the remainder of the down time

Run Status

  • Accelerator wrapped up maintenance and resumed beam around 1930 yesterday. Wien angle was changed
  • All contaminated materials removed from target enclosure yesterday
  • Hall went back to restricted access this morning to continue with the configuration change
    • Bogdan and Jack began compass field measurements around the target enclosure around 2pm
    • HCal move progressing
    • Target cell is removed
    • Possible status update from Jesse
  • Last update: A final RF module needs replaced in replaced BPM VME crate. The work will need to be validated by Simona.
    • Possible status update from Simona
  • Other Subsystem updates?

Future Plans

  • Hall in restricted access for remainder of configuration change
    • Hall access is open to subsystem experts after 1700 and before 0800
    • Please check with RC for access beyond DAQ bunker during the day

Other Topics

Experimental progress (for MCC planning meeting): Based on UPDATED Run Coordinator instructions (new this week)

  • 6.1% of accumulated charge for PRODUCTION data (13.55 Coulombs out of 232)
  • 28% Experimental run time (39 out of 141 days)

Current Run Plan

Configuration change.

Need updated 'standard' run plan for 3-pass data taking

NMR measurement Restore beam

Run Plan 1.) He-3 runs, at 45uA of beam current. Use ps3==0, ps5=12, all others -1. Take four runs that should each last for one hour. 2.) Move the target to the pickup coil position, and the TO should run an NMR sweep to measure the polarization. 3.) Return to step 1.)

  • Keep an eye out for drifting BPM positions
  • Log all DAQ crashes in the excel worksheet along with the reason for the crash
  • If beam has been down due to accelerator issues perform a checkout of the beam profile by doing harp scans and check beam centering on target using the Carbon hole target, see Simona's procedure: [1], If in doubt about whether you should go through the beam checkout procedure, call this number: 7578693131.
  • Cross check the rate with what is expected: [2]
  • Cross check with the expected ion chamber readings: [3]

Notices in Effect for Shift Crew

Current Issues

  • Given the delay and unfeasibility of completing the configuration change tomorrow, when can we expect to complete the configuration change?
  • Shift Openings: Wednesday OWL, Thursday OWL. [4]

Planned Accesses to the Hall

  • In Restricted Access. Configuration change ongoing.

Special Requests